# Facial Expression Analysis and Its Potentials in IoT Systems: A Contemporary Survey

### v1

@@ 4.2.1. Ensemble learning
Ensemble learning can enhance the accuracy and robustness of dynamic MaE recognition by frame aggregation that can combine temporal dynamic features across multiple frames and frame-level emotion classification results In deep dynamic MaE recognition, the frame aggregation can be implemented at decision-level (Kahou et al., 2013, 2016) and feature-level (Nguyen et al., 2022; Liu et al., 2022b, 2023b; Pan et al., 2024).
@@ 4.2.2. Explicit spatio-temporal learning
Explicit spatio-temporal learning focuses on constructing deep models that are designed to extract spatio-temporal information from the dynamic MaE datasets. The explicit spatio-temporal learning methods can incorporate the temporal information into the encoded features by using a sequence of frames within a sliding window. The recurrent neural network (RNN) (Elman, 1990) and its advanced version (i.e., long short-term memory, LSTM) (Hochreiter and Schmidhuber, 1997) have been used in dynamic MaE recognition due to the ability to process sequential data (Zhang et al., 2017; Lee et al., 2020; Wang et al., 2022b; Uddin et al., 2021; ZhouXuan, 2021; Pei et al., 2021; Yu et al., 2018b; An and Liu, 2020; Abdullah et al., 2020). For instance, Zhang et al. (Zhang et al., 2017) employed an RNN to articulate morphological changes and dynamic evolution of MaEs by exploiting key facial regions based on facial landmarks. Zhi et al. (Zhi and Wan, 2019) proposed a lightweight LSTM-based method with less complexity to produce a sparse representation for dynamic MaE recognition. Lee et al. (Lee et al., 2020) introduced attention-guided convolutional LSTM to capture dynamic spatio-temporal information. By leveraging the depth and thermal sequences as guidance priors, the proposed model can guide the model to focus on discriminative visual regions. An LSTM autoencoder was used to learn temporal dynamics for dynamic MaE recognition (Uddin et al., 2021). Besides, the 3D deep models (e.g., 3D CNN (Ji et al., 2012)) have also been adopted to extract spatio-temporal features directly for dynamic MaE recognition (Hasani and Mahoor, 2017; Deng et al., 2019; Cai et al., 2018). Hasani et al. (Hasani and Mahoor, 2017) combined 3D Inception-ResNets (Szegedy et al., 2017) with an LSTM unit to capture spatio-temporal information. Deng et al. (Deng et al., 2019) proposed a 3D CNN framework consisting of a stem layer, 3D Inception-ResNets structure, and gated recurrent unit (GRU) layer to process video data. To further improve the performance of the model, the island loss (Cai et al., 2018) was incorporated to increase inter-class differences while minimizing intra-class variations. Khanna et al. (Khanna et al., 2024) developed an end-to-end spatio-temporal deep model that integrates residual networks (He et al., 2016) and DenseNet (Huang et al., 2017) for MaE recognition.
@@ 4.2.3. Multi-task learning
Multi-task learning in deep dynamic MaE recognition focuses on capturing the dynamic variation patterns of MaEs over time in order to enhance the understanding of expression changes in video clips. By fusing local dynamic features extracted from a part-based module, Yu et al. (Yu et al., 2020b) proposed a multi-task module that can capture the subtle variations of MaEs from both local and global features. Jin et al. (Jin et al., 2021) utilized a multi-task framework involving AU detection and MaE recognition to pre-train the visual feature extractor. Then, both visual and audio features from videos were concatenated and fed into a transformer encoder to extract temporal features for recognition. Xie et al. (Xie et al., 2023) expanded the scope of auxiliary tasks by incorporating dynamic MaE recognition from multiple modalities, such as text, audio, and video.
@@ 4.2.4. Attention-based learning
Attention-based learning in deep dynamic MaE recognition focuses on both spatial and temporal information. Compared with attention modules in deep static MaE recognition, attention modules in deep dynamic MaE recognition operate across the temporal and even channel dimensions to extract richer video features. Furthermore, these attention modules can integrate multimodal features to derive more comprehensive spatio-temporal representations.
@@
Moreover, transformers have been leveraged to integrate multi-modal features for dynamic MaE recognition. For instance, Zhang et al. (Zhang et al., 2022) proposed a transformer-based framework that fuses static visual features with dynamic multi-modal information derived from text, audio, and video. In the proposed framework of (Zhang et al., 2022), static frame features act as a guiding mechanism for the multi-modal fusion process to ensure the alignment of temporal dynamics with spatial cues from individual frames. Building on the framework in (Zhang et al., 2022), Zhang et al. (Zhang et al., 2024) introduced an advanced multi-modal transformer framework with improved methods for feature fusion and extraction. Specifically, the framework employs three transformer-based encoders, each dedicated to extracting features from a specific modality—visual, audio, or text—while enabling modality-specific adaptation. Unlike the earlier framework (Zhang et al., 2022), which focused on integrating multi-modal data using static frame guidance, the subsequent method (Zhang et al., 2024) achieves more robust multi-modal feature integration by aligning semantic information across modalities. This is accomplished by mapping features from all modalities into a shared latent space anchored in visual features, thereby supporting a more unified representation of multi-modal data.
@@ 4.2.5. Self-supervised learning
Self-supervised learning focuses on capturing changes in expressions across spatial and temporal dimensions in the deep dynamic MaE recognition. Moreover, self-supervised models can learn spatio-temporal features from unlabeled video data. Since the self-supervised learning models can interpret videos by learning the order of frames in human action recognition (Fernando et al., 2017; Sharma et al., 2019), Song et al. (Song et al., 2021) introduced a self-supervised framework with a rank loss mechanism to generate dynamic MaE features. By sorting preceding and following frames based on their distance from a central frame, the proposed framework can generate robust dynamic MaE features. Sun et al. (Sun et al., 2023b) proposed a large-scale self-supervised model for pre-training a dynamic MaE recognition system using extensive unlabeled facial video data. Chumachenko et al. (Chumachenko et al., 2024) proposed a multi-modal approach that combined complementary features from diverse modalities. The multi-modal approach employed two self-supervised learning encoders pre-trained on audio sequences and static images, and then fine-tuned using audio-visual dynamic videos.
4.2.6. Potential limitations of different methods on deep dynamic MaE recognition.
Ensemble learning in dynamic MaE recognition increases both computational cost and system complexity. For example, feature-level frame aggregation requires substantial computation due to the large volume of frame-level features and complex network structures e.g., 3D CNNs, attention mechanisms, or transformers. Explicit spatiotemporal learning methods, such as 3D CNNs and RNN variants such as LSTM, GRU, are highly computationally expensive during training and inference, which may be impractical for resource-constrained applications. Furthermore, these models for the explicit spatiotemporal learning method with large amounts of parameters, are prone to overfitting during the training stage, especially when the MaE data is limited and the MaE changes are subtle within video sequences. The selection of appropriate auxiliary tasks in multi-task learning is a major challenge for MaE recognition. Irrelevant tasks or auxiliary tasks with significantly different levels of difficulty can disrupt model training. In multi-modal and multi-task dynamic recognition models, integrating data from text, audio, and video requires precise alignment and synchronization. Misalignment between these modalities can degrade the model’s performance, particularly when capturing subtle expression changes. Attention mechanisms can effectively integrate facial region information and capture long-range dependences, but they can also introduce latency issues, particularly in models like transformers, which suffer from increased computational and memory requirements as the input sequence length grows. Although self-supervised learning captures sequential information, it struggles to model long-term correlations, especially when MaE changes are not immediately obvious.
@@ 5.1. MiE Spotting
MiE spotting, referring to locating temporal segments in a specific video clip, is an important step for holographic MiE analysis. The key description of MiE in a video clip is to identify three time spots: (1) the onset refers to the initial frame where the emotion of an MiE becomes discernible; (2) the apex frame is the frame within the MiE clip that exhibits the highest intensity of emotion; and (3) the offset frame marks the end of the emotion. The current works on MiE spotting can be categorized into apex-frame spotting and interval spotting.
@@ 5.1.1. Apex frame spotting
The apex frame reflects the highest intensity of an MiE within a video clip and implies the ground-truth emotion of individuals. Traditional apex frame spotting methods exploited the algorithms based on handcrafted features, such as the descriptors of LBP and its variants due to their effectiveness and robustness. Pfister et al. (Pfister et al., 2011b) pioneered a MiE spotting approach for short spontaneous expressions. The proposed method in (Pfister et al., 2011b) leverages local binary patterns from three orthogonal planes (LBP-TOP) (Zhao and Pietikainen, 2007) to achieve notable results compared to manual detection methods. Yan et al. (Yan et al., 2015) proposed to utilize two handcrafted descriptors (i.e., the constrained local model (Cristinacce et al., 2006) and LBP) to detect the ROI of each face and to extract subsequent features for MiE analysis. To address the limitations of LBP-TOP in capturing essential information, Esmaeili et al. (Esmaeili and Shahdi, 2020) proposed Cubic-LBP—an improved version of LBP. Compared with LBP, Cubic-LBP can extract fine-grained features from 15 planes and integrate the obtained histograms. Although the use of additional planes could enrich the features, Cubic-LBP also introduced redundant data that increases computational expenditure. To alleviate the computational expenditure, Esmaeili et al. (Esmaeili et al., 2022) proposed an intelligent Cubic-LBP by leveraging CNNs to select relevant planes.
@@ 5.1.2. Interval spotting
Interval spotting refers to detecting the onset and offset in a long MiE video with various constraints, such as, MaE, head movement, and blinks. While more challenging than apex frame spotting, interval spotting has broader practical applications. For interval spotting, traditional methods typically set thresholds to locate key points by comparing feature differences across entire videos. For example, Li et al. (Li et al., 2017b) proposed a method based on comparison of difference characteristics. Specifically, the proposed method identified LBP as a more effective baseline technique compared to the optical flow histogram. By incorporating the multi-scale analysis and a sliding window, Tran et al. (Tran et al., 2017) scaled the video sequence and obtained corresponding samples for spotting MiE intervals. Optical flow has been used in interval spotting due to its ability to describe motion information. Patel et al. (Patel et al., 2015) introduced a heuristic algorithm with optical flow for MiE interval spotting. The algorithm emphasized the continuity of motion direction. By analyzing the direction of motion vectors over time, it ensured that detected movements extracted by optical flow were consistent.
Wang et al. (Wang et al., 2016, 2017) introduced the main directional maximal difference (MDMD) to characterize facial motion by calculating the largest magnitude difference in the main direction of optical flow. The proposed method was later adopted as a baseline in the third facial MiE grand challenge (Li et al., 2020b). Han et al. (Han et al., 2018) enhanced feature difference analysis by combining LBP with the main directional mean optical flow (MDMO) (Liu et al., 2015), which leveraged both texture and motion features for complementary insights. Guo et al. (Guo et al., 2021) extended this work by incorporating both magnitude and angle information from optical flow to detect local MiE movements.
@@
MiE interval spotting methods focusing on handcrafted features and optical flow mentioned above require appropriate threshold settings, which is susceptible to the selected features and unexpected facial motion. To address the challenges, deep model approaches have emerged as promising alternatives. To distinguish the MaE and MiE in a long video sequence, Pan et al. (Pan et al., 2020) presented a local bilinear CNN to detect the fine-grained facial regions associated with MiE. Takalkar et al. (Takalkar et al., 2021) developed LGAttNet, a dual-attention framework that integrated local and global facial features for improved spotting. Gu et al. (Gu et al., 2023) proposed a lightweight deep model to predict the likelihood of a frame being part of a MiE interval for overfitting mitigation. To capture spatio-temporal features, Wang et al. (Wang et al., 2021) designed a 2D spatial and 1D temporal convolutional model. In addition, many studies (Yu et al., 2021a; Yap et al., 2022) utilized 3D CNNs to extract robust spatio-temporal features. Zhou et al. (Zhou et al., 2022) further enhanced performance by integrating 3D CNNs with bidirectional encoder representations from transformers (BERTs) for onset-offset detection. Moreover, approaches leveraging AU analysis (Yin et al., 2023; Yang et al., 2023) can incorporate prior knowledge and reduce noise interference to improve MiE spotting performance.
5.1.3. Potential limitations of different methods on MiE spotting.
Traditional handcrafted descriptors are robust and offer good interoperability. For apex frame spotting, methods like LBP-TOP can handle low-intensity MiEs effectively. However, handcrafted descriptors are typically limited to local information and may fail to fully capture global dynamic patterns. CNN-based sliding window approaches can locate apex frames effectively, but are less time-efficient for long videos. The performance of sliding window methods is sensitive to the choice of window size and step size, impacting both detection speed and accuracy. Apex frame spotting also requires high-quality data, which may not always be available in practical applications.
Traditional interval spotting methods using handcrafted descriptors rely on threshold-based detection to locate key points by comparing feature differences across entire videos. This makes them highly sensitive to feature selection and unexpected facial movements. The fine-tuning of thresholds is a challenge in diverse real-world scenarios. Deep learning models, although successful in MiE interval spotting, require large, labeled datasets. Given that MiEs are subtle and brief, models can overfit to the specific characteristics of the training data.

### v3

@@ 4.2.1. Ensemble learning
Ensemble learning can enhance the accuracy and robustness of dynamic MaE recognition by frame aggregation that can combine temporal dynamic features across multiple frames and frame-level emotion classification results. In deep dynamic MaE recognition, the frame aggregation can be implemented at decision-level (Kahou et al., 2013, 2016) and feature-level (Nguyen et al., 2022; Liu et al., 2022b, 2023b; Pan et al., 2024).
@@ 4.2.2. Explicit spatio-temporal learning
Explicit spatio-temporal learning focuses on constructing deep models that are designed to extract spatio-temporal information from the dynamic MaE datasets. The explicit spatio-temporal learning methods can incorporate the temporal information into the encoded features by using a sequence of frames within a sliding window. The recurrent neural network (RNN) (Elman, 1990) and its advanced version (i.e., long short-term memory, LSTM) (Hochreiter and Schmidhuber, 1997) have been used in dynamic MaE recognition due to the ability to process sequential data (Zhang et al., 2017; Lee et al., 2020; Wang et al., 2022d; Uddin et al., 2021; ZhouXuan, 2021; Pei et al., 2021; Yu et al., 2018b; An and Liu, 2020; Abdullah et al., 2020). For instance, Zhang et al. (Zhang et al., 2017) employed an RNN to articulate morphological changes and dynamic evolution of MaEs by exploiting key facial regions based on facial landmarks. Zhi et al. (Zhi and Wan, 2019) proposed a lightweight LSTM-based method with less complexity to produce a sparse representation for dynamic MaE recognition. Lee et al. (Lee et al., 2020) introduced attention-guided convolutional LSTM to capture dynamic spatio-temporal information. By leveraging the depth and thermal sequences as guidance priors, the proposed model can guide the model to focus on discriminative visual regions. An LSTM autoencoder was used to learn temporal dynamics for dynamic MaE recognition (Uddin et al., 2021). Besides, the 3D deep models (e.g., 3D CNN (Ji et al., 2012)) have also been adopted to extract spatio-temporal features directly for dynamic MaE recognition (Hasani and Mahoor, 2017; Deng et al., 2019; Cai et al., 2018). Hasani et al. (Hasani and Mahoor, 2017) combined 3D Inception-ResNets (Szegedy et al., 2017) with an LSTM unit to capture spatio-temporal information. Deng et al. (Deng et al., 2019) proposed a 3D CNN framework consisting of a stem layer, 3D Inception-ResNets structure, and gated recurrent unit (GRU) layer to process video data. To further improve the performance of the model, the island loss (Cai et al., 2018) was incorporated to increase inter-class differences while minimizing intra-class variations. Khanna et al. (Khanna et al., 2024) developed an end-to-end spatio-temporal deep model that integrates residual networks (He et al., 2016) and DenseNet (Huang et al., 2017) for MaE recognition.
@@ 4.2.3. Multi-task learning
Multi-task learning in deep dynamic MaE recognition focuses on capturing the dynamic variation patterns of MaEs over time in order to enhance the understanding of expression changes in video clips. By fusing local dynamic features extracted from a part-based module, Yu et al. (Yu et al., 2020) proposed a multi-task module that can capture the subtle variations of MaEs from both local and global features. Jin et al. (Jin et al., 2021) utilized a multi-task framework involving AU detection and MaE recognition to pre-train the visual feature extractor. Then, both visual and audio features from videos were concatenated and fed into a transformer encoder to extract temporal features for recognition. Xie et al. (Xie et al., 2023) expanded the scope of auxiliary tasks by incorporating dynamic MaE recognition from multiple modalities, such as text, audio, and video.
@@ 4.2.4. Attention-based learning
Attention-based learning in deep dynamic MaE recognition focuses on both spatial and temporal information. Compared with its deep static variants, attention modules in deep dynamic MaE recognition operate across the temporal and even channel dimensions to extract richer video features. Furthermore, the attention modules can integrate multimodal features to derive more comprehensive spatio-temporal representations.
@@
Moreover, transformers have been leveraged to integrate multi-modal features for dynamic MaE recognition. For instance, Zhang et al. (Zhang et al., 2022) proposed a transformer-based framework that fuses static visual features with dynamic multi-modal information derived from text, audio, and video. In the proposed framework of (Zhang et al., 2022), static frame features act as a guiding mechanism for the multi-modal fusion process to ensure the alignment of temporal dynamics with spatial cues from individual frames. Building on the framework in (Zhang et al., 2022), Zhang et al. (Zhang et al., 2024) introduced an advanced multi-modal transformer framework with improved methods for feature fusion and extraction. Specifically, the framework employs three transformer-based encoders, each dedicated to extracting features from a specific modality—visual, audio, or text—while enabling modality-specific adaptation. Unlike the earlier framework (Zhang et al., 2022) that focused on integrating multi-modal data using static frame guidance, the subsequent method (Zhang et al., 2024) achieves more robust multi-modal feature integration by aligning semantic information across modalities by mapping features from all modalities into a shared latent space anchored in visual features, thereby supporting a more unified representation of multi-modal data.
@@ 4.2.5. Self-supervised learning
Self-supervised learning focuses on capturing changes in expressions across spatial and temporal dimensions in the deep dynamic MaE recognition. Moreover, self-supervised models can learn spatio-temporal features from unlabeled video data. Since the self-supervised learning models can interpret videos by learning the order of frames in human action recognition (Fernando et al., 2017; Sharma et al., 2019), Song et al. (Song et al., 2021) introduced a self-supervised framework with a rank loss mechanism to generate dynamic MaE features. By sorting preceding and following frames based on their distance from a central frame, the proposed framework can generate robust dynamic MaE features. Sun et al. (Sun et al., 2023a) proposed a large-scale self-supervised model for pre-training a dynamic MaE recognition system using extensive unlabeled facial video data. Chumachenko et al. (Chumachenko et al., 2024) proposed a multi-modal approach that combined complementary features from diverse modalities. The multi-modal approach employed two self-supervised learning encoders pre-trained on audio sequences and static images, and then fine-tuned using audio-visual dynamic videos.
4.2.6. Integrated Comparisons & Practical Insights on Deep Dynamic MaE Recognition
Deep dynamic MaE recognition enables models to capture the temporal evolution of emotions by expanding facial expression analysis to video sequences. Various learning paradigms have been proposed to extract spatiotemporal features critical for accurate dynamic MaE recognition. Recent studies explore a range of strategies, including frame-level integration (ensemble learning (Kahou et al., 2013, 2016; Nguyen et al., 2022; Liu et al., 2022b, 2023b; Pan et al., 2024)), modeling spatio-temporal dynamics (explicit spatio-temporal learning (Elman, 1990; Hochreiter and Schmidhuber, 1997; Zhang et al., 2017; Lee et al., 2020; Wang et al., 2022d; Uddin et al., 2021; ZhouXuan, 2021; Pei et al., 2021; Yu et al., 2018b; An and Liu, 2020; Abdullah et al., 2020; Zhi and Wan, 2019; Ji et al., 2012; Hasani and Mahoor, 2017; Deng et al., 2019; Cai et al., 2018; Szegedy et al., 2017; Khanna et al., 2024; He et al., 2016; Huang et al., 2017), attention‑based learning (Meng et al., 2019; Liu et al., 2020a; Sun et al., 2021; Xia et al., 2022; Chen et al., 2023a; Zhao and Liu, 2021; Huang et al., 2021; Ma et al., 2022; Zhang et al., 2022; Zhao et al., 2023b; Wang et al., 2024; Chen et al., 2024; Zhang et al., 2024)), and fusing multi-source features (multi-task learning (Yu et al., 2020; Jin et al., 2021; Xie et al., 2023)) for improved dynamic MaE recognition. However, these approaches still face challenges such as the lack of explicit temporal dependency modeling, high training complexity, and the scarcity of well-aligned benchmark datasets. Self-supervised learning (Fernando et al., 2017; Sharma et al., 2019; Song et al., 2021; Sun et al., 2023a; Chumachenko et al., 2024) partially mitigates the data‑scarcity issue by exploiting unlabeled videos, but the cost of lengthy training cycles and substantial compute requirements limits its practical deployment. A detailed comparison of deep dynamic MaE recognition methods is provided in Appendix 3.2.
The temporal nature of dynamic MaE recognition introduces unique bottlenecks. For example, traditional 3D CNNs that process entire video clips for analysis can increased latency, which is a critical issue for edge devices handling real-time streams. One effective solution is to reduce the complexity of the input representation. For example, (Deng et al., 2019) leverage facial landmarks to extract temporal MaE cues while integrating spatial appearance features from individual frames. Similarly, (Zhao et al., 2023b) demonstrate that expressive spatiotemporal dynamics can be captured using only landmark sequences, bypassing raw pixel input. In addition, (Song et al., 2021) further show that dynamic representations can even be derived from static images, supporting the feasibility of key-frame-based recognition pipelines. An alternative solution focuses on architectural efficiency. Recent approaches (Wang et al., 2024; Chen et al., 2024) adopt lightweight backbones such as ResNet-18 for spatial encoding, coupled with transformer-based temporal modules to capture global temporal dependencies. This modular design enables flexible control over the number of input frames, providing a practical trade-off between latency and recognition performance.
Another major challenge in dynamic MaE recognition is the scarcity of large-scale labeled datasets. Existing datasets used for practical applications(e.g., DFEW, AFEW) are limited in size, which hinders the the training of deep temporal models with strong generalization capabilities. A widely adopted solution to this problem is the use of self-supervised pre-training on large-scale unlabeled but domain-relevant data. This approach enables models to learn generalized MaE representations, which can be fine-tuned for downstream dynamic facial expression recognition tasks. For instance, (Sun et al., 2023a; Chumachenko et al., 2024) demonstrate that applying a masked autoencoding reconstruction objective on unlabeled MaE data can enhance performance on dynamic recognition tasks. In addition, (Song et al., 2021) show that learning temporal ordering between adjacent frames provides valuable temporal supervision, allowing the model to capture dynamic dependencies and improve its ability to model subtle emotional transitions across time.
@@ 5.1. MiE Spotting
MiE spotting, referring to locating temporal segments in a specific video clip, is an important step for holographic MiE analysis. The key description of MiE in a video clip is to identify three time spots: (1) the onset refers to the initial frame where the emotion of a MiE becomes discernible; (2) the apex frame is the frame within the MiE clip that exhibits the highest intensity of emotion; and (3) the offset frame marks the end of the emotion. The current works on MiE spotting can be categorized into apex-frame spotting and interval spotting.
@@ 5.1.1. Apex frame spotting
The apex frame reflects the highest intensity of a MiE within a video clip and implies the ground-truth emotion of individuals. Traditional apex frame spotting methods exploited the algorithms based on handcrafted features, such as the descriptors of LBP and its variants due to their effectiveness and robustness. Pfister et al. (Pfister et al., 2011b) pioneered a MiE spotting approach for short spontaneous expressions. The proposed method in (Pfister et al., 2011b) leverages local binary patterns from three orthogonal planes (LBP-TOP) (Zhao and Pietikainen, 2007) to achieve notable results compared to manual detection methods. Yan et al. (Yan et al., 2015) proposed to utilize two handcrafted descriptors (i.e., the constrained local model (Cristinacce et al., 2006) and LBP) to detect the ROI of each face and to extract subsequent features for MiE analysis. To address the limitations of LBP-TOP in capturing essential information, Esmaeili et al. (Esmaeili and Shahdi, 2020) proposed Cubic-LBP—an improved version of LBP. Compared with LBP, Cubic-LBP can extract fine-grained features from 15 planes and integrate the obtained histograms. Although the use of additional planes could enrich the features, Cubic-LBP also introduced redundant data that increases computational expenditure. To alleviate the computational expenditure, Esmaeili et al. (Esmaeili et al., 2022) proposed an intelligent Cubic-LBP by leveraging CNNs to select relevant planes.
@@ 5.1.2. Interval spotting
Interval spotting refers to detecting the onset and offset in a long MiE video with various constraints, such as, MaE, head movement, and blinks. While more challenging than apex frame spotting, interval spotting has broader practical applications. For interval spotting, traditional methods typically set thresholds to locate key points by comparing feature differences across entire videos. For example, Li et al. (Li et al., 2017b) proposed a method based on comparison of difference characteristics. Specifically, the proposed method identified LBP as a more effective baseline technique compared to the optical flow histogram. By incorporating the multi-scale analysis and a sliding window, Tran et al. (Tran et al., 2017) scaled the video sequence and obtained corresponding samples for spotting MiE intervals. Optical flow has been used in interval spotting due to its ability to describe motion information. Patel et al. (Patel et al., 2015) introduced a heuristic algorithm with optical flow for MiE interval spotting. The algorithm emphasized the continuity of motion direction. By analyzing the direction of motion vectors over time, it ensured that detected movements extracted by optical flow were consistent. Wang et al. (Wang et al., 2016, 2017) introduced the main directional maximal difference (MDMD) to characterize facial motion by calculating the largest magnitude difference in the main direction of optical flow. The proposed method was later adopted as a baseline in the third facial MiE grand challenge (Li et al., 2020b). Han et al. (Han et al., 2018) enhanced feature difference analysis by combining LBP with the main directional mean optical flow (MDMO) (Liu et al., 2015), which leveraged both texture and motion features for complementary insights. Guo et al. (Guo et al., 2021) extended this work by incorporating both magnitude and angle information from optical flow to detect local MiE movements.
@@
MiE interval spotting methods focusing on handcrafted features and optical flow mentioned above require appropriate threshold settings, which is susceptible to the selected features and unexpected facial motion. To address the challenges, deep model approaches have emerged as promising alternatives. To distinguish the MaE and MiE in a long video sequence, Pan et al. (Pan et al., 2020) presented a local bilinear CNN to detect the fine-grained facial regions associated with MiE. Takalkar et al. (Takalkar et al., 2021) developed LGAttNet, a dual-attention framework that integrated local and global facial features for improved spotting. Gu et al. (Gu et al., 2023) proposed a lightweight deep model to predict the likelihood of a frame being part of a MiE interval for overfitting mitigation. To capture spatio-temporal features, Wang et al. (Wang et al., 2021) designed a 2D spatial and 1D temporal convolutional model. In addition, many studies (Yu et al., 2021b; Yap et al., 2022) utilized 3D CNNs to extract robust spatio-temporal features. Zhou et al. (Zhou et al., 2022) further enhanced performance by integrating 3D CNNs with bidirectional encoder representations from transformers (BERTs) for onset-offset detection. Moreover, approaches leveraging AU analysis (Yin et al., 2023; Yang et al., 2023) can incorporate prior knowledge and reduce noise interference to improve MiE spotting.
5.1.3. Integrated Comparisons & Practical Insights on MiE spotting
MiE spotting is a precursor to successful MiE recognition, which demands precise temporal sensitivity to localize transient facial cues often embedded within long, expression-neutral sequences. Early methods relied heavily on handcrafted motion descriptors, such as optical flow (Patel et al., 2015; Wang et al., 2016, 2017) and LBP on temporal planes (Pfister et al., 2011b; Esmaeili and Shahdi, 2020), while more recent approaches have adopted deep learning techniques that encode temporal patterns from full video sequences or sparse frames (Zhang et al., 2018a; Yee et al., 2022; Pan et al., 2020; Takalkar et al., 2021; Gu et al., 2023; Wang et al., 2021; Yu et al., 2021b; Yap et al., 2022; Zhou et al., 2022; Yin et al., 2023; Yang et al., 2023). These methods attempt to balance the trade-off between computational cost, robustness to noise, and the fidelity of motion representation. Despite advancements, spotting remains an underexplored and error-prone stage due to ambiguous expression boundaries and individual-specific motion variances. A detailed comparison of deep dynamic MaE recognition methods is provided in Appendix 4.1.
Despite a decade of progress, reliable MiE spotting remains elusive in real‑world settings. Three interrelated issues are highlighted below. Lack of a unified evaluation protocol. Current studies report results under heterogeneous conditions, which hinders fair comparison. For apex spotting, some authors process isolated MiE clips whereas others search for apex frames in long, mixed‑expression videos, leading to incomparable recall and precision numbers. Interval‑level spotting has likewise evolved from early “enlarged-window” hit tests (Li et al., 2017b) to the now popular Intersection‑over‑Union (IoU $\geq$ 0.5) criterion (Pan et al., 2020). A consensus benchmark is urgently needed: (1) apex spotting should be evaluated in long, unconstrained sequences that include distractor expressions; and (2) interval spotting should adopt IoU‑based metrics, which align with modern object‑detection practice. Sensitivity to illumination and head motion. Hand‑crafted descriptors encode pixel‑wise differences and optical flow, making them brittle under pose changes and lighting fluctuations. Li et al. (Li et al., 2017b), for instance, report frequent false positives caused by eye blinks. Moreover, their pipeline requires a temporal interpolation model that alone costs $\approx$22 s per video, pushing total inference time beyond usability for online applications. Limited recall in deep models for long videos. Deep architectures overcome the temporal interpolation model bottleneck by feeding fixed‑stride windows or variable‑length clips directly to 3‑D CNNs, temporal transformers, or graph–temporal hybrids. Although this reduces preprocessing latency and boosts interval‑level F1 scores (e.g., 38.34% on CAS(ME)2 and 37.28% on SAMM‑LV (Yin et al., 2023)), performance drops when focusing exclusively on MiE (15.38% and 19.84%, respectively). These results indicate that current spotting models are still insufficiently sensitive to the sparse, low‑intensity cues that distinguish MiEs from surrounding MaEs.